A family has two cars. The first car has a fuel efficiency of 40 miles per gallon of gas and the second has a fuel efficiency of 15 miles per gallon of gas. During one particular week, the two cars went a combined total of 1600 miles, for a total gas consumption of 65 gallons. How many gallons were consumed by each of the two cars that week?
Answer by ikleyn(52786)

.

Let x = the number of gallons consumed by the 1-st car and y = that for the 2-nd car.


Then you have this system of 2 equations in 2 unknowns


      x +   y =   65       (1)    (65 gallons of gas, total)

    40x + 15y = 1600       (2)    (miles, total)


There are many methods to solve the system.
For example, you may use the Elimination method.
For it, multiply eq(1) by 40 (both sides).  Keep eq(2) as is


    40x + 40y = 65*40      (1')   

    40x + 15y = 1600       (2')    


Next, subtract eq(2') from eq(1).  You will get


          40y - 15y = 65*40 - 1600 

          25y        = 1000

            y        = 1000/25 = 40.


ANSWER.  2-nd car consumed 40 gallons of gas;  1-st car consumed  the rest  65-40 = 25 gallons.

Solved.

--------------------

To learn on how to algebreze and to solve similar problems, see the lessons
    - Solution of the linear system of two equations in two unknowns by the Substitution method
    - Solution of the linear system of two equations in two unknowns by the Elimination method
    - Solution of the linear system of two equations in two unknowns using determinant
    - Geometric interpretation of the linear system of two equations in two unknowns
    - Useful tricks when solving systems of 2 equations in 2 unknowns by the Substitution method
    - Solving word problems using linear systems of two equations in two unknowns

    - Oranges and grapefruits
    - Using systems of equations to solve problems on tickets
    - Three methods for solving standard (typical) problems on tickets
    - Using systems of equations to solve problems on shares
    - Using systems of equations to solve problems on investment
    - Two mechanics work on a car
    - The Robinson family and the Sanders family each used their sprinklers last summer
    - Roses and vilolets
    - Counting calories and grams of fat in combined food
    - A theater group made appearances in two cities
    - HOW TO algebreze and solve this problem on 2 equations in 2 unknowns
    - One unusual problem to solve using system of two equations
    - Solving mentally word problems on two equations in two unknowns
    - OVERVIEW of lessons on solving systems of two linear equations in two unknowns

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.
in this site.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.